# I3-MRec: Invariant Learning with Information Bottleneck for Incomplete Modality Recommendation

### v1

@@ 5.1.1. Datasets
We evaluate our method on three widely used real-world datasets: Amazon Baby, Amazon Clothing, and TikTok, following the 5-core setting adopted in prior works (Zhou et al., 2023a; Wei et al., 2020; Ong and Khong, 2025). The Amazon datasets [^6] http://jmcauley.ucsd.edu/data/amazon. consist of user-item interactions derived from product ratings. For each item, we extract 4096-dimensional visual features using VGG16 (Simonyan and Zisserman, 2015) (from the second fully connected layer), and 1024-dimensional textual features using BERT (Devlin et al., 2019), based on the concatenation of brand, title, description, and category, following (Liu et al., 2023). The TikTok dataset (Bai et al., 2024) contains user viewing histories on short videos, with pre-extracted visual, acoustic, and textual features. We directly use the released modality features and data splits to ensure consistency and fair comparison. Following (Liu et al., 2023), we adopt the same training, validation, and test partition strategy. Dataset statistics are summarized in Table 1.
@@ 5.1.3. Baselines
To verify the efficacy of our proposed method, we conduct comprehensive benchmarking against a diverse set of state-of-the-art (SOTA) recommender models, which can be broadly categorized into three main groups: Unimodal Recommendation Methods: MF-BPR  (Rendle et al., 2009), InvPref (Wang et al., 2022), and LightGCN (He et al., 2020). These methods rely solely on user-item interaction data to infer user preferences, without incorporating auxiliary content features. Multimodal Recommendation Methods: feature-based models (VBPR (He and McAuley, 2016), BM3 (Zhou et al., 2023a), InRL (Du et al., 2022)), graph-based models (GRCN (Wei et al., 2020), MMGCN (Wei et al., 2019), LATTICE (Zhang et al., 2021), DualGNN (Wang et al., 2021)), and hybrid models (MGCN (Yu et al., 2023), SMORE (Ong and Khong, 2025)). These methods incorporate multimodal item features to enhance recommendation accuracy. Incomplete Modality Recommendation Methods: These models are specifically designed to alleviate modality-missing problems, which include representative approaches (MILK (Bai et al., 2024), SIBRAR (Ganhör et al., 2024)), and Generative models(MoDiCF (Li et al., 2025)).
@@ 5.3. Improvements on Modality Missingness
In Figure  3, we evaluate the performance of the proposed model and MRS (e.g., VBPR, LATTICE, SMORE) across different types of missing modalities. To simulate the real-world scenario, we select the MTMT strategy as the experimental setting [^8] The MTMT setting simulates realistic and diverse missing cases, including mismatches between training and test missing modalities (e.g., training with missing image, testing with missing text or both. From the observation of the figure, we can find that 1) I3-MRec consistently outperformed all baselines across all modality-missing scenarios. This demonstrates the effectiveness of our framework in learning robust representations even in the absence of two modalities. 2) Benefiting from the introduction of invariant learning and information bottleneck, I3-MRec can yield compact yet effective representation, which maximally preserves user preference information while minimizing dependence on the raw modality content. As shown in the figure, our model maintains stable performance regardless of which modality is missing. while traditional MRS models are especially sensitive to the absence of visual features, leading to a sharp performance drop. 3) Beyond robustness under modality-missing conditions, I3-MRec also performs competitively in full-modality settings, achieving results on par with the recent work SMORE. Interestingly, the motivation of SMORE is to capture both uni-modal and fusion preferences while simultaneously suppressing modality noise. This suggests that reducing modality redundancy through the Information Bottleneck principle not only benefits generalization in modality-missing scenarios but also contributes to better modeling of user preferences in the complete modality scenario.
Figure 4. (a) Recall@20 under varying missing modality rates on the Amazon Baby dataset. (b) Normalized performance with respect to the full-modality setting: Under the 90% modality missing rate, the proposed method preserves 88% of its original performance, with a degradation of only 12%.
@@
Interestingly, MoDiCF performs well under low missing ratios, retaining 98% of full-modality performance at 10%. However, as missingness increases (50%-90%), its effectiveness declines sharply, with recall dropping by nearly 20%, revealing the limitations of generation-based strategies in modality-missing scenarios. In contrast, I3-MRec learns compact and robust representations, maintaining strong performance even when most modalities are unavailable. These results demonstrate the model’s robustness and generalization under various modality-missing conditions.
@@ 5.4. Ablation Study
To further examine the role of the IB module, we introduce I3-MRec, a variant that incorporates both IRM and IB but omits the information compression (minimization) term. Its performance lags behind the full model, particularly under the MTMT setting, indicating that without information compression, the model becomes more susceptible to noisy or missing modality inputs during training.
@@ 6. Conclusion
In this work, we make a clear and principled contribution to robust multimodal recommendation by explicitly identifying two fundamental properties essential for effective recommendation under missing modality conditions: (i) cross-modal feature invariance and (ii) compact, preference-oriented representation. Grounded in these properties, we propose I3-MRec, a novel and generalizable framework that integrates Invariant Risk Minimization (IRM) and the Information Bottleneck (IB) principle. Specifically, IRM enables the model to learn stable user preference representations across modality-specific environments, while IB encourages the extraction of essential, task-relevant information and reduces reliance on noisy or modality missingness. This design allows I3-MRec to achieve robust and accurate recommendations in realistic modality-missing scenarios. Extensive experiments on three real-world datasets demonstrate that I3-MRec consistently outperforms unimodal, multimodal, and modality-missing aware baselines across a range of missing ratios and evaluation settings (FTMT and MTMT). These results validate the robustness and generalization capabilities of our framework.

### v2

@@ 5.1.1. Datasets
We evaluate our method on three widely used real-world datasets: Amazon Baby, Amazon Clothing, and TikTok, following the 5-core setting adopted in prior works (Zhou et al., 2023a; Wei et al., 2020; Ong and Khong, 2025). The Amazon datasets [^7] http://jmcauley.ucsd.edu/data/amazon. consist of user-item interactions derived from product ratings. For each item, we extract 4096-dimensional visual features using VGG16 (Simonyan and Zisserman, 2015) (from the second fully connected layer), and 1024-dimensional textual features using BERT (Devlin et al., 2019), based on the concatenation of brand, title, description, and category, following (Liu et al., 2023a). The TikTok dataset (Bai et al., 2024) contains user viewing histories on short videos, with pre-extracted visual, acoustic, and textual features. We directly use the released modality features and data splits to ensure consistency and fair comparison. Following (Liu et al., 2023a), we adopt the same training, validation, and test partition strategy. Dataset statistics are summarized in Table 1.
@@ 5.1.3. Baselines
To verify the efficacy of our proposed method, we conduct comprehensive benchmarking against a diverse set of state-of-the-art (SOTA) recommender models, which can be broadly categorized into three main groups: Unimodal Recommendation Methods: MF-BPR  (Rendle et al., 2009), InvPref (Wang et al., 2022), and LightGCN (He et al., 2020). These methods rely solely on user-item interaction data to infer user preferences, without incorporating auxiliary content features. Multimodal Recommendation Methods: feature-based models (VBPR (He and McAuley, 2016), BM3 (Zhou et al., 2023a), InRL (Du et al., 2022)), graph-based models (GRCN (Wei et al., 2020), MMGCN (Wei et al., 2019), LATTICE (Zhang et al., 2021), DualGNN (Wang et al., 2021)), and hybrid models (MGCN (Yu et al., 2023), SMORE (Ong and Khong, 2025)). Incomplete Modality Recommendation Methods: These models are specifically designed to alleviate modality-missing problems, which include representative approaches (MILK (Bai et al., 2024), SIBRAR (Ganhör et al., 2024)), and Generative models(MoDiCF (Li et al., 2025)).
@@ 5.3. Improvements on Modality Missingness
In Figure  3, we evaluate the performance of the proposed model and MRS (e.g., VBPR, LATTICE, SMORE) across different types of missing modalities. To simulate the real-world scenario, we select the MTMT strategy as the experimental setting [^9] The MTMT setting simulates realistic and diverse missing cases, including mismatches between training and test missing modalities (e.g., training with missing image, testing with missing text or both. From the observation of the figure, we can find that 1) I3-MRec consistently outperformed all baselines across all modality-missing scenarios. This demonstrates the effectiveness of our framework in learning robust representations even in the absence of two modalities. 2) Benefiting from the introduction of invariant learning and information bottleneck, I3-MRec can yield compact yet effective representation, which maximally preserves user preference information while minimizing dependence on the raw modality content. As shown in the figure, our model maintains stable performance regardless of which modality is missing. while traditional MRS models are especially sensitive to the absence of visual features, leading to a sharp performance drop. 3) Beyond robustness under modality-missing conditions, I3-MRec also performs competitively in full-modality settings, achieving results on par with the recent work SMORE. Interestingly, the motivation of SMORE is to capture both uni-modal and fusion preferences while simultaneously suppressing modality noise. This suggests that reducing modality dependence through the Information Bottleneck principle benefits generalization in modality-missing scenarios.
Figure 4. (a) Recall@20 under varying missing modality rates on the Amazon Baby dataset. (b) Normalized performance with respect to the full-modality setting.
@@
Interestingly, MoDiCF performs well under low missing ratios, retaining 98% of full-modality performance at 10%. However, as missingness increases (50%-90%), its effectiveness declines sharply, with recall dropping by nearly 20%, revealing the limitations of generation-based strategies in modality-missing scenarios. In contrast, I3-MRec learns compact and effective representations, maintaining stable performance even when most modalities are unavailable, demonstrating the model’s robustness and generalization.
@@ 5.4. Ablation Study
To further examine the role of the IB module, we introduce I3-MRec, a variant that incorporates both IRM and IB but omits the information compression (minimization) term. Its performance lags behind the full model, particularly under the MTMT setting, indicating that without information compression, the model becomes more susceptible to missing modality inputs during training.
@@ 6. Conclusion
In this work, we make a principled contribution to robust multimodal recommendation by explicitly identifying two fundamental properties essential for effective recommendation under missing modality conditions: (i) cross-modal preferences invariance and (ii) compact, preference-oriented representation. Grounded in these properties, we propose I3-MRec, a novel and generalizable framework that integrates Invariant Risk Minimization (IRM) and the Information Bottleneck (IB) principle. Specifically, IRM enables the model to learn stable, reference-oriented representations across modality-specific environments, while IB encourages the model to preserve preference-relevant information and reduce reliance on modality-specific information. This design allows I3-MRec to achieve robust and accurate recommendations in realistic modality-missing scenarios. Extensive experiments on three real-world datasets demonstrate that I3-MRec consistently outperforms unimodal, multimodal, and modality-missing aware baselines across a range of missing ratios and evaluation settings (FTMT and MTMT). These results validate the robustness and generalization capabilities of our framework.